OpenStudy (anonymous):

consider the plane -x+4y+2z+6=0 a) determine a normal vector,n to the plane b) determine the coordinates of two points ,s and t, on the plane c)determine st d)show that st is perpendicular to n

OpenStudy (anonymous):

please help

OpenStudy (turingtest):

take the gradient of the equation for the plane do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

do you know how to do cross products?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

...that will make this difficult do you know how to take a partial derivative?

OpenStudy (anonymous):

dervative i think so

OpenStudy (turingtest):

do you know how to find the determinant of a matrix?

OpenStudy (anonymous):

no havent studied that

OpenStudy (anonymous):

its calculus 12

OpenStudy (turingtest):

ok, I'm gonna try to do this the only way I know how what is the partial derivative of the function for the plane with respect to x ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it -1+4+2=0

OpenStudy (anonymous):

?

OpenStudy (turingtest):

in vector form, yes -i+4j+2k

OpenStudy (turingtest):

that is the gradient, so you apparently do know how to take it after all

OpenStudy (turingtest):

you realise that the partial with respect to x is just -1 you did \(f_x\hat i+f_y\hat j+f_z\hat k\)

OpenStudy (turingtest):

the gradient of a function is\[\nabla f=f_x\hat i+f_y\hat j+f_z\hat k\]

OpenStudy (turingtest):

that is a vector that is perpendicular to the surface so the vector you want is\[\nabla f(x,y,z)=\vec n=\langle-1,4,2\rangle\]

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

so how do i determine the coordinates of s and t on plane

OpenStudy (turingtest):

well, unless you did not provide some info, you just need to pick any two points in the plane you can do that by just plugging in numbers for example, let x=0, y=0, then you can figure out what z is do that and get two points

OpenStudy (anonymous):

aha but the quwstion doesnt have any graphs

OpenStudy (turingtest):

so just pick two random points in the plane

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

so for t i do the same?

OpenStudy (turingtest):

yes, we need to get the coordinates of any two points so we can get a vector in the plane we are then going to prove that that vector is perpendicular to \(\vec n\) for part c)

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

so point s is (1,2,-6.5)

OpenStudy (anonymous):

t is (2,3,-8)

OpenStudy (turingtest):

if x=1 and y=2 I get z=-0.5 try to pick whole numbers though if you can x=0, y=0 are obvious choices; or letting any two variables be zero, then finding the third

OpenStudy (anonymous):

oh k thx

OpenStudy (turingtest):

the (2,3,-8) works just fine though, don't get me wrong

OpenStudy (anonymous):

yea nvm

OpenStudy (turingtest):

ah you were right about the other point so you want to use s is (1,2,-6.5) and t is (2,3,-8) ??? can you form a vector between them?

OpenStudy (anonymous):

yea st= t-s i guess

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

i got (1,1,-14,5)

OpenStudy (anonymous):

-14.5

OpenStudy (turingtest):

-8-(-6.5)=-1.5

OpenStudy (anonymous):

so how do we find st perpendicular to n

OpenStudy (anonymous):

oh my bad

OpenStudy (turingtest):

what do you know about the dot product of two perpendicular vectors?

OpenStudy (anonymous):

d.m=0 so d is perpendicular to n

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

so we use st to find d?

OpenStudy (turingtest):

the last part just asks us to confirm that the vector st it perpendicular to \(\vec n\) so dot them, and if it comes out to be zero we have proved that they are perpendicular

OpenStudy (anonymous):

so in this case st.n=0

OpenStudy (turingtest):

it should, let's check

OpenStudy (turingtest):

\[\vec{st}\cdot\vec n=\langle1,1,-1.5\rangle\cdot\langle-1,4,2\rangle\]

OpenStudy (anonymous):

i got [-1,4,-3]

OpenStudy (turingtest):

careful, dot products are scalar\[\vec{st}\cdot\vec n=\langle1,1,-1.5\rangle\cdot\langle-1,4,2\rangle=-1+4-3=0\]

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

it matchs

OpenStudy (turingtest):

cool how even those random points you chose prove our theorem it always matches, or else math fails :)

OpenStudy (anonymous):

yea i think so thx for your help very much

OpenStudy (turingtest):

welcome :)